A beam of initially unpolarized light passes throu
Answer #1


Intensity oflight coming from 1st polarizer I1 = I0/2

Intensity oflight coming from 2nd polarizer I2 = I1*(cos21.7)^2 = (I0/2)*(cos21.7)^2

Intensity oflight coming from 3rd polarizer I3 = I2*(cos52.9)^2 = (I0/2)*(cos21.7)^2*(cos52.9)^2


I3/I0 = 0.157 <<-------------answer